# AnyRIR: Robust Non-intrusive Room Impulse Response Estimation in the Wild

### v1

@@ Abstract
I Introduction
@@
However, estimating RIRs in uncontrolled real-world environments remains a challenge [1]. Conventional methods, such as exponential sine sweeps, require controlled and often intrusive setups that are impractical in public or dynamic spaces. Moreover, non-stationary noise sources (e.g., footsteps or speech) corrupt these measurements [4, 5, 6]. Inspired by the widespread presence of background music in public venues and the availability of clean references through music identification algorithms such as those used by Shazam [7] (first described in [8]), we propose a robust and non-intrusive method for estimating RIRs in the wild using music as the excitation signal. Fig. 1 illustrates how our method can estimate an RIR from music recorded in a cafe.
While stationary noise with a Gaussian distribution is less detrimental, non-stationary noise severely degrades deconvolution-based estimation [9]. MOSAIC [6] introduces a way to remove non-stationary noise from exponential sine sweep measurements using median filtering, but it requires at least three repeated measurements with exactly the same excitation signal, which makes it somewhat inflexible. As an alternative, we employ a regression-based approach using the $\ell_{1}$-norm approximation, which is known for its robustness to outliers [10]. This makes it particularly suitable for handling non-stationary noise, since it suppresses the influence of transient disturbances in the measurement.
Deconvolution using the $\ell_{1}$-norm has been studied in various domains. In underwater acoustics, it has been applied to both the noise and the filter, using the Alternating Direction Method of Multipliers (ADMM) in the time domain to promote sparsity in the filter [11]. In contrast, this work develops a method for deconvolution in the time–frequency domain without imposing sparsity on the filter. Similarly, in image processing, $\ell_{1}$-norm minimization has been used to suppress impulsive noise during deconvolution [12, 13].
@@
Figure 1: (Top) A noisy environment, such as cafe, where music is playing and is reverberating, and (bottom) the AnyRIR principle of estimating the RIR with the help of music recognition.
II Problem Formulation
@@
Figure 2: The proposed AnyRIR method estimates the RIR by minimizing the time-frequency residual between the recorded music and the clean music convolved with the previous RIR estimate. The main tools used are the IRLS and LSMR methods.
@@
Given an RIR $\bm{h}$ and an excitation signal $\bm{x}$, the measured signal $\bm{y}$ at a microphone can be modeled as $\bm{y}=\bm{x}\mathbin{\ast}\bm{h}+\bm{n},$ where $\mathbin{\ast}$ denotes the convolution operation and $\bm{n}$ represents additive background noise. Let $\bm{x}$ have length $L$, and $\bm{h}$ have length $N$, then $\bm{y}$ has length M=L+N1M=L+N-1. We aim to solve the inverse problem of estimating the unknown RIR $\bm{h}$ from a known pair of excitation and measurement signals, $\bm{x}$ and $\bm{y}$. When the noise $\bm{n}$ is assumed to be Gaussian, modeling stationary background noise, the deconvolution problem can be approached with the $\ell_{2}$-norm loss. However, real-world noise often contains non-stationary impulsive components, which is better modeled by heavy-tailed, non-Gaussian distributions. Such impulsive components act as outliers in the residual and degrade the performance of $\ell_{2}$-based estimators. To enhance robustness against outliers, we adopt the $\ell_{1}$-norm loss [10] and write the convolution as a multiplication operator by forming the Toeplitz matrix $\bm{X}$ of $\bm{x}(t)$:
@@
III Proposed Estimation Method
The proposed method, AnyRIR, solves the deconvolution problem using the $\ell_{1}$-norm minimization (2) in the time-frequency domain, where $\bm{X}$ and $\bm{S}$ are large matrices. This leads to a time and memory limitation when using standard optimization libraries, such as CVXPY [16]. There are various efficient alternatives, such as ADMM [11]. In particular, we choose the Iterative Reweighted Least Squares (IRLS) method along with a matrix-free iterative method called the Least-Squares Minimal Residual (LSMR) to handle the least-squares subproblem [17], as shown in Fig. 2, while avoiding to form $\bm{X}$ and $\bm{S}$ explicitly. The IRLS algorithm is formulated as a sequence of weighted least squares problems:
@@ III Proposed Estimation Method
For small-scale problems where $\bm{X}$ is well-conditioned, the least squares subproblem can be solved analytically using the normal equations. However, in our setting, where the excitation signal is very long, forming the normal equation becomes memory-intensive. In addition, the convolutional structure of $\bm{X}$ often leads to ill-conditioned systems [18], resulting in numerical instability. Therefore, we employ an iterative solver, such as the Conjugate Gradient method [19] and LSQR [20]. In particular, we use LSMR, which is a matrix-free method for sparse and or fast linear operator, offering improved robustness when handling ill-conditioned matrices [17].
@@
Figure 3: Number of LSMR iterations required for convergence with and without EQ preconditioning.
@@
IV Evaluation
@@
IV-A Dataset
To simulate music played in a typical public environment, we convolve music signals with RIRs and add both stationary and non-stationary noise. Specifically, we generated four different songs using Suno AI[22], covering styles characteristic of café environments: pop, melodic acoustic instrumentation, coffee-shop ambiance, male and female vocals, light percussion, and full-band arrangements. For non-stationary noise, we use the AID dataset [23], which contains anechoic recordings of real-world interferers such as footsteps or coughs. RIRs are drawn from the MIT Acoustical Reverberation Scene Statistics Survey [24], which contains 271 RIR representing everyday acoustic environments.
IV-B Presence of Non-stationary Noise
Figure 4: (Left) Music signal convolved with an RIR with added stationary and non-stationary noise: $\bm{y}$. (Center) Original non-stationary noise. (Right) Weights $\bm{w}$ computed by AnyRIR to suppress the bins containing the non-stationary noise.
Figure 5: EDCs from different deconvolution methods compared to the ground truth, showing the superiority of AnyRIR.
@@ IV-B Presence of Non-stationary Noise
As an upper bound, we consider performance on signals without non-stationary noise. Stationary noise remains challenging to suppress, but its effect can be averaged out by extending measurement signals, e.g., to 2–3 min. Table I shows that AnyRIR performs equivalently to the standard $\ell_{2}$ method in the absence of non-stationary noise, while providing robustness when such noise is present. Results are reported as mean and standard deviation across the 50 examples.
TABLE I: Comparison of RIR estimation error ($\bm{h}$ error in dB, mean $\pm$ std) under different noise conditions. AnyRIR maintains low error in both cases, while baseline methods degrade substantially in the presence of non-stationary noise.
@@ IV-C Music Codec
Figure 6: Effect of codec mismatch. When the measurement and excitation signals use different codecs, the estimated RIR exhibits increased noise in the late tail, resulting in $\sim$15 dB higher error compared to the same-codec condition.
IV-D Real-world Evaluation
@@
Figure 7: EDC estimated from a real-world recording.
V Conclusions

### v2

@@ Abstract
Index Terms—  Acoustic measurements, deconvolution, music.
1 Introduction
@@
However, estimating RIRs in uncontrolled real-world environments remains a challenge [11]. Conventional methods, such as exponential sine sweeps, require controlled and often intrusive setups that are impractical in public or dynamic spaces. Moreover, non-stationary noise sources (e.g., footsteps or speech) corrupt these measurements [5, 14, 15]. Inspired by the widespread presence of background music in public venues and the availability of clean references through music identification algorithms such as Shazam [shazam2025, wang_industrial-strength_2003], we propose a robust and non-intrusive method for estimating RIRs in the wild using music as the excitation signal. Fig. 1 illustrates how our method can estimate an RIR from music recorded in a cafe.
@@
Fig. 1: (Top) A noisy environment, such as cafe, where music is playing and is reverberating, and (bottom) usage of AnyRIR to estimate the RIR with the help of existing music recognition algorithm.
While stationary noise with a Gaussian distribution can often be handled adequately by $\ell_{2}$-based deconvolution methods, non-stationary noise violates the underlying assumptions and severely degrades estimation performance [17]. MOSAIC [15] introduces a way to remove non-stationary noise from exponential sine sweep measurements using median filtering, but it requires at least three repeated measurements with exactly the same excitation signal, which makes it somewhat inflexible. As an alternative, we adopt a regression-based approach using the $\ell_{1}$-norm loss, which is known for its robustness to outliers [Boyd:2004jj]. This makes it particularly suitable for handling non-stationary noise, since it suppresses the influence of transient disturbances in the measurement. Although $\ell_{1}$-norm regularization has previously been used to promote sparsity in RIR estimation [crocco2015room], inducing sparsity in the filter is not our objective. Instead, the $\ell_{1}$-norm is used solely as the data fidelity term to improve robustness to non-stationary noise, without imposing sparsity or other constraints on the RIR.
@@
Fig. 2: The proposed AnyRIR method estimates the RIR by minimizing the time-frequency residual between the recorded music and the clean music convolved with the previous RIR estimate. The main tools used are the IRLS and LSMR methods.
Robust deconvolution using the $\ell_{1}$-norm has been explored in other domains. In underwater acoustics, it has been applied to channel estimation under impulsive noise using time-domain ADMM [2]. In image processing, $\ell_{1}$-norm minimization has been widely used to suppress impulsive noise during deconvolution [1, 4]. In contrast, our approach operates in the time–frequency domain and does not impose sparsity on the filter, focusing instead on robustness to non-stationary noise in real-world acoustic measurements.
2 Problem Formulation
@@
Given an RIR $\bm{h}$ and an excitation signal $\bm{x}$, the measured signal $\bm{y}$ at a microphone can be modeled as $\bm{y}=\bm{x}\mathbin{\ast}\bm{h}+\bm{n},$ where $\mathbin{\ast}$ denotes the convolution operation and $\bm{n}$ represents additive background noise. Let $\bm{x}$ have length $L$, and $\bm{h}$ have length $N$, then $\bm{y}$ has length M=L+N1M=L+N-1. We aim to solve the inverse problem of estimating the unknown RIR $\bm{h}$ from a known pair of excitation and measurement signals, $\bm{x}$ and $\bm{y}$. When the noise $\bm{n}$ is assumed to be Gaussian, modeling stationary background noise, the deconvolution problem can be approached with the $\ell_{2}$-norm loss. However, real-world noise often contains non-stationary impulsive components, which is better modeled by heavy-tailed, non-Gaussian distributions. Such impulsive components act as outliers in the residual and degrade the performance of $\ell_{2}$-based estimators. To enhance robustness against outliers, we adopt the $\ell_{1}$-norm loss [Boyd:2004jj] and write the convolution as a multiplication operator by forming the Toeplitz matrix $\bm{X}$ of $\bm{x}(t)$:
@@
3 Proposed Estimation Method
The proposed method, AnyRIR, solves the deconvolution problem using the $\ell_{1}$-norm minimization (2) in the time-frequency domain, where $\bm{X}$ and $\bm{S}$ are large matrices. This leads to a time and memory limitation when using standard optimization libraries, such as CVXPY [diamond2016cvxpy]. There are various efficient alternatives, such as ADMM [2]. In particular, we choose the Iterative Reweighted Least Squares (IRLS) method along with a matrix-free iterative method called the Least-Squares Minimal Residual (LSMR) to handle the least-squares subproblem [7], as shown in Fig. 2, while avoiding to form $\bm{X}$ and $\bm{S}$ explicitly. The IRLS algorithm is formulated as a sequence of weighted least squares problems:
@@ 3 Proposed Estimation Method
For small-scale problems where $\bm{X}$ is well-conditioned, the least squares subproblem can be solved analytically using the normal equations. However, in our setting, where the excitation signal is very long, forming the normal equation becomes memory-intensive. In addition, the convolutional structure of $\bm{X}$ often leads to ill-conditioned systems [3], resulting in numerical instability. Therefore, we employ an iterative solver, such as the Conjugate Gradient method [Shewchuk.1994] and LSQR [12]. In particular, we use LSMR, which is a matrix-free method for sparse and or fast linear operator, offering improved robustness when handling ill-conditioned matrices [7].
@@
Fig. 3: Number of LSMR iterations required for convergence with and without EQ preconditioning.
4 Evaluation
@@
4.1 Dataset
To simulate music played in a typical public environment, we convolve music signals with RIRs and add both stationary and non-stationary noise. Specifically, we generated four different songs using Suno AI[suno2025], covering styles characteristic of café environments: pop, melodic acoustic instrumentation, coffee-shop ambiance, male and female vocals, light percussion, and full-band arrangements. For non-stationary noise, we use the AID dataset [8], which contains anechoic recordings of real-world interferers such as footsteps or coughs. RIRs are from the MIT Acoustical Reverberation Scene Statistics Survey [18], which contains 271 RIR representing everyday acoustic environments.
4.2 Presence of Non-stationary Noise
Fig. 4: (Left) Music signal convolved with an RIR with added stationary and non-stationary noise: $\bm{y}$. (Center) Original non-stationary noise. (Right) Weights $\bm{w}$ computed by AnyRIR to suppress the bins containing the non-stationary noise.
Fig. 5: EDCs from different deconvolution methods compared to the ground truth, showing the superiority of AnyRIR.
@@ 4.2 Presence of Non-stationary Noise
As an upper bound, we consider performance on signals without non-stationary noise. Stationary noise remains challenging to suppress, but its effect can be averaged out by extending measurement signals, e.g., to 2–3 min. Table 1 shows that AnyRIR performs equivalently to the standard $\ell_{2}$ method in the absence of non-stationary noise, while providing robustness when such noise is present. Results are reported as mean and standard deviation across the 50 examples.
Table 1: Comparison of RIR estimation error ($\bm{h}$ error in dB, mean $\pm$ std) under different noise conditions. AnyRIR maintains low error in both cases, while baseline methods degrade substantially in the presence of non-stationary noise.
@@ 4.3 Music Codec
Fig. 6: Effect of codec mismatch. When the measurement and excitation signals use different codecs, the estimated RIR exhibits increased noise in the late tail, resulting in $\sim$15 dB higher error compared to the same-codec condition.
4.4 Real-world Evaluation
@@
Fig. 7: EDC estimated from a real-world recording.
5 Conclusions